OpenStudy (anonymous):

find the diameter of a sphere that has a surface area of 100(3.14)

OpenStudy (ddcamp):

\[SA_{sphere} = 4 \pi r^2 \\ 100 = 4 (3.14) r^2\]

OpenStudy (ddcamp):

\[r^2 = (\frac{ d }{ 2 })^2 = \frac{ d^2 }{ 4 }\]

OpenStudy (anonymous):

\[S=4 \pi r^2\] \[r = \sqrt{S/4\pi}\] \[d = 2 \sqrt{S/4 \pi}\] \[d = 2 \sqrt{25\pi}\] \[d = 10\sqrt{\pi}\]